OpenStudy (anonymous):

[4_07]Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (anonymous):

Can you write an equation for your new line in point-slope form?

OpenStudy (anonymous):

Im not sure what point-slope is lol I think its slope intercept form though...

OpenStudy (anonymous):

Point-slope form is the easiest way to write an equation from a point and a slope. It looks like this: \[y-y _{1}=m(x-x_1)\] where m is the slope and \[(x_1,y_1)\] is the point.

OpenStudy (anonymous):

Now Im confused. . . Im just working on Alg1 so I dont really know

OpenStudy (anonymous):

That's OK. Here's an example of an equation in point-slope form: \[y-1 = 3(x-2)\] The slope of the line is 3, and the line contains the point (1,2).

OpenStudy (anonymous):

Oh, no I dont think it can be that bc all the answers are just y without another number. ..

OpenStudy (compassionate):

Hello, my name is Stephen and I will walk you through the steps. Remember, you're finding the parallel line. Parallel lines have the exact same slope. I can represent this in a photo.  The only difference in a parallel line is the y-value. So, we know it passes through the line y = 3x - 6 We know the slope stays the same. The slope is 3 We are also given two points (2, -4) Slope: 3 Points (2, -4) Now that that information and plug it into the equation\[y - y1 = m(x - x1) \rightarrow y - (-4) = 3(x - 2)\] Now you ned to simplify the equation. \[y - (-4) = 3(x - 2) = y + 4 = 3x - 6 \rightarrow \rightarrow y = 3x - 6 -4 \] Now simplify the final part of the equation.

OpenStudy (anonymous):

y=3x-10?

OpenStudy (compassionate):

Correct. Good job! Come back to OpenStudy if you have any other questions

OpenStudy (anonymous):

OH lol thanks! c:

OpenStudy (compassionate):

It is always encouraged to give best answers.